There are 4 different mathematics books and 5 different science books in how many ways can the books be arranged on a shelf if?

Answer by ikleyn(52777)

.

If WHAT ?


Without any "if", 4 + 5 = 9 different items can be arranged in 9! = 9*8*7*6*5*4*3*2*1 = 362880 ways.     ANSWER


It is the number of PERMUTATIONS of 9 different items.


--------------------------

On Permutations,  see introductory lessons
    - Introduction to Permutations
    - PROOF of the formula on the number of Permutations
    - Problems on Permutations

    - OVERVIEW of lessons on Permutations and Combinations
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.